OpenStudy (anonymous):

Graph the conic section. 16y2 – 25x2 = 400

OpenStudy (anonymous):

@lucus1697 @terenzreignz @telijahmed @agent0smith @christianfilms11 @hartnn

OpenStudy (anonymous):

25x^2 – 16y^2 = 400 25x^2/400 – 16y^2/400 = 400/400 x^2/16 – y^2/25 = 1

OpenStudy (anonymous):

@freshkid944

OpenStudy (anonymous):

im still confused how do you graph it which one do you graph? @christianfilms11

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

thank so much